OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (solomonzelman):

Let's subtract the second equation from the first equation. -2x + y = 1 - -4x + y = -1 --------------------------------- -2x - - 4x y - y = 1 - - 1 -2x+4x + 0 = 1+1 -2x=2 x=-1

OpenStudy (solomonzelman):

Can you solve for y?

OpenStudy (anonymous):

I don't really understand how. Can you solve it & show me all of the steps again ?

OpenStudy (solomonzelman):

What you are doing is subtract the first equation from the second ok?

OpenStudy (anonymous):

Okay so 1 - (-1) = 2.

OpenStudy (solomonzelman):

yes 1-(-1) = 1+1 = 2 correct. now lets subtract the variable y in the second equation from variable y in the first equation, Ok?

OpenStudy (solomonzelman):

y-y =0 correct? so so far we have __x - 0 = 2

OpenStudy (anonymous):

But now what..

OpenStudy (solomonzelman):

subtract the variable x in the second equation from the variable x in the first equation. lets write the subtraction out, -2x - (-4x)

OpenStudy (anonymous):

I think the answer is (-1,-3) .. Is that right ?

OpenStudy (solomonzelman):

can we first solve the equation, before jumping to the answer?